# Adaptive Neural Networks for Intelligent Data-Driven Development

### v1

@@ 1 Introduction
Some of the main challenges in developing ML-based systems include identifying critical variables that affect model performance and establishing robust performance guarantees for various scenarios. To address these challenges, researchers have developed different methodologies for safely building systems with ML components[3, 21]. In these approaches, strong emphasis is placed on ensuring data coverage for a well-defined operational domain design (ODD), followed by carefully orchestrated model development and testing processes. However, even with thorough dataset curation, unanticipated scenarios, objects, or conditions can arise once the vehicle is operational in the real world. These out-of-distribution (OoD) instances pose a significant risk to performance, as most ML models exhibit a notable drop in accuracy when encountering inputs with distributions that deviate from their training data [18].
@@
2 Methodology
@@ 2.1 Overview
2.2 Initial Segmentation Network
Figure 3: Adaptive neural network architecture with and OoD detection mechanism
@@ 2.2 Initial Segmentation Network
where $C$ is the number of components per GMM, $\pi_{yc}$ is the component mixture weight for component $c$ of class $y$, $\mu_{yc}$,$\Sigma_{yc}$ are the mean and covariance matrix respectively, and $\mathcal{N}$ is the Gaussian distribution. The parameters of the decoder $\theta$ are optimized by minimizing the cross-entropy loss on the training set $\mathcal{D}_{in}$ over the posterior output of the GMM $\{\phi_{y}\}_{y=1}^{|Y|}$:
@@ 2.3 OoD Detection
To address this issue, we propose a likelihood ratio-based formulation that adapts to evolving in-distribution classes without requiring explicit re-training of the OoD detection module. Similar to [20], our method uses the likelihood ratio formulation for detecting OoD objects, but we adjust the inlier score to appropriately adjust to newly incorporated classes, preventing them from being detected as OoD once they have been learned.
@@
2.4 Retrieval
To enable effective search within a large-scale unlabeled dataset, we propose a retrieval framework that allows users to find relevant samples based on either semantic descriptions or visual references. Different retrieval needs depend on the user’s use case. When a user is searching for instances based on general object categories—such as “trailer” or “e-scooter”—a text-based retrieval approach similar to [19] is ideal. In such cases, joint vision-language embeddings enable semantic retrieval, allowing users to discover relevant objects even when exact matches are unavailable. Conversely, when a user has a specific reference object and wants to retrieve visually similar instances—such as different viewpoints of the same anomaly—an image-based retrieval method typically provides a more precise search mechanism based purely on visual similarity.
@@ 2.5 Continual Learning
$\begin{split}\mathcal{L}&=\underbrace{-\frac{1}{N}\sum_{i=1}^{N}\log p(y_{i}|x_{i},G_{\text{new}})}_{\text{inlier loss}}\\ &-\underbrace{\frac{\lambda}{N}\sum_{i=1}^{N}\log\frac{p(x_{i}|y_{i},G_{\text{new}})}{\sum_{y^{\prime}\in\mathcal{Y}}p(x_{i}|y^{\prime},G_{\text{old}})}}_{\text{contrastive loss}}\end{split}$ (7)
@@
3 Experiments
3.1 Experimental Setup
@@ 3.1.1 Metrics
3.2 Results
@@
Table 1: Comparison of OoD detection methods across SOS and WOS, and the inlier performance on Cityscapes
3.2.1 Inlier and OoD Detection Performance
Table 1 compares our approach with two recent state-of-the-art methods (UEM [12] and RbA [11]) across both datasets. Averaged across both datasets, our method performs 0.05 less than RbA and 3.8 less than UEM in the F1 scores. However, our method requires significantly fewer parameters: 3.2 times fewer than UEM and 6.4 times fewer than RbA. This reduction in model size makes our approach more suitable for resource-constrained environments, such as in-car systems, where available memory and computational capacity are limited. Moreover, the efficiency of our model allows for faster inference, making it a viable choice for real-time applications where low latency is critical.
3.2.2 Retrieval Performance
@@
Table 2: Retrieval Performance on SOS and WOS
Table 2 shows the retrieval performance of the image-based and text-based retrieval on the SOS and WOS datasets. Averaged across the two datasets, the image-based retrieval performs better than the text-based retrieval by 4.75 mp, while having $\approx 7$ times less parameters. However, the image-based retrieval requires a clustering step to obtain the centroids, which the user can select as queries, in this examples we used simple $k$-means clustering on the embeddings to obtain the clusters, and the parameter $k$ was obtained using the elbow method. However, if the parameter $k$ is set to a value which is lower than the number of objects in the dataset, then some of the potential object queries might be missed. If the number of objects cannot be roughly estimated before clustering, then text-based retrieval might be a better option.
@@
3.2.3 Continual Learning
Table 3 shows the performance on each incrementally learned class and the original 19 classes. All retrieved instances are split into 80:20 splits for training and testing. The mIoU on the validation set of the newly learned class is less than the mIoU for the original inlier class with a margin of 2.5 mIoU points. This decrease in performance is most likely due to the nature of the newly learned objects being more difficult to learn due to the lack of sample points. Additionally, we find that our model performs better on larger objects for the original and incrementally added classes than on smaller objects. This discrepancy can be attributed to the network’s architecture, which only upsamples feature maps to 2/7 of the original image resolution, with the final output bilinearly upsampled by a factor of 3.5. This leads to the loss of fine details, making it more challenging for the model to classify smaller objects accurately.
@@ 3.2.3 Continual Learning
Table 3: IoU results for the original 19 classes and the incrementally learned classes.
4 Conclusion and Future Work
@@
Acknowledgment. The research leading to these results is funded by the German Federal Ministry for Economic Affairs and Climate Action within the project “just better DATA”.

### v3

@@ I INTRODUCTION
Some of the main challenges in developing ML-based systems include identifying critical variables that affect model performance and establishing robust performance guarantees for various scenarios. To address these challenges, researchers have developed different methodologies for safely building systems with ML components[4, 5]. In these approaches, strong emphasis is placed on ensuring data coverage for a well-defined operational domain design (ODD), followed by carefully orchestrated model development and testing processes. However, even with well-curated datasets, real-world deployment can expose models to unanticipated, out-of-distribution (OoD) inputs that degrade the performance of machine learning algorithms, a critical concern in safety-sensitive domains like autonomous driving [6].
@@
II SMART DATA LOOP FOR LEARNING NEW CLASSES
@@ II-A Overview
II-B Initial Segmentation Network
Figure 3: Adaptive neural network architecture with an OoD detection mechanism.
@@ II-B Initial Segmentation Network
where $C$ is the number of components per GMM, $\pi_{yc}$ is the component mixture weight for component $c$ of class $y$, $\mu_{yc}$,$\Sigma_{yc}$ are the mean and covariance matrix respectively, and $\mathcal{N}$ is the Gaussian distribution. The parameters of the decoder $\theta$ are optimized by minimizing the cross-entropy loss on the training set $\mathcal{D}_{in}$ over the posterior output of the parameters of the GMM $\phi$:
@@ II-C OoD Detection
To address this issue, we propose a likelihood ratio-based formulation that adapts to evolving in-distribution classes without requiring explicit re-training of the OoD detection module. Inspired by Nayal et al. [13], our method ensures that the OoD score appropriately adjusts to newly incorporated classes, preventing them from being detected as OoD once they have been learned.
@@
II-D Retrieval
To enable effective search within a large-scale unlabeled dataset, we propose a retrieval framework that allows users to find relevant samples based on either semantic descriptions or visual references. Different retrieval needs depend on the user’s use case. When a user is searching for instances based on general object categories—such as “trailer” or “e-scooter”—a text-based retrieval approach is ideal. In such cases, joint vision-language embeddings enable semantic retrieval, allowing users to discover relevant objects even when exact matches are unavailable [14]. Conversely, when a user has a specific reference object and wants to retrieve visually similar instances—such as different viewpoints of the same anomaly—an image-based retrieval method typically provides a more precise search mechanism based purely on visual similarity.
@@ II-E Continual Learning
$\begin{split}\mathcal{L}&=\underbrace{-\frac{1}{N}\sum_{i=1}^{N}\log p(y_{i}|x_{i},\phi_{\text{new}})}_{\text{inlier loss}}\\ &-\underbrace{\frac{\lambda}{N}\sum_{i=1}^{N}\log\frac{p(x_{i}|y_{i},\phi_{\text{new}})}{\sum_{y^{\prime}\in\mathcal{Y}}p(x_{i}|y^{\prime},\phi_{\text{old}})}}_{\text{contrastive loss}}\end{split}$ (7)
@@
III EVALUATIONS
III-A Experimental Setup
@@ III-A1 Metrics
III-B Results
TABLE I: Comparison of OoD detection methods across SOS and WOS, and the inlier performance on Cityscapes
@@ III-B Results
III-B1 Inlier and OoD Detection Performance
Table I compares our approach with two recent state-of-the-art methods (UEM [13] and RbA [12]) across both datasets. Averaged across both datasets, our method performs 0.05 less than RbA and 3.8 less than UEM in the F1 scores. However, our method requires significantly fewer parameters: 3.2 times fewer than UEM and 6.4 times fewer than RbA. This reduction in model size makes our approach more suitable for resource-constrained environments, such as in-car systems, where available memory and computational capacity are limited. Moreover, the efficiency of our model allows for faster inference, making it a viable choice for real-time applications where low latency is critical.
III-B2 Retrieval Performance
TABLE II: Retrieval Performance on SOS and WOS
@@ III-B2 Retrieval Performance
Table II shows the retrieval performance of the image-based and text-based retrieval on the SOS and WOS datasets. Averaged across the two datasets, the image-based retrieval performs better than the text-based retrieval by 4.75 mp, while having $\approx 7$ times less parameters. However, the image-based retrieval requires a clustering step to obtain the centroids, which the user can select as queries, in this examples we used simple $k$-means clustering on the embeddings to obtain the clusters, and the parameter $k$ was obtained using the elbow method. However, if the parameter $k$ is set to a value which is lower than the number of objects in the dataset, then some of the potential object queries might be missed. If the number of objects cannot be roughly estimated before clustering, then text-based retrieval might be a better option.
@@
III-B3 Continual Learning
Table III shows the performance on each incrementally learned class and the original 19 classes. All retrieved instances are split into 80:20 splits for training and testing. The mIoU on the validation set of the newly learned class is less than the mIoU for the original inlier class with a margin of 2.5 mIoU points. This decrease in performance is most likely due to the nature of the newly learned objects being more difficult to learn due to the lack of sample points. Additionally, we find that our model performs better on larger objects for the original and incrementally added classes than on smaller objects. This discrepancy can be attributed to the network’s architecture, which only upsamples feature maps to 2/7 of the original image resolution, with the final output bilinearly upsampled by a factor of 3.5. This leads to the loss of fine details, making it more challenging for the model to classify smaller objects accurately.
@@ III-B3 Continual Learning
TABLE III: IoU results for the original 19 classes and the incrementally learned classes.
@@
IV CONCLUSIONS AND FUTURE WORK